# A new paradigm of mortality modeling via individual vitality dynamics

### v2

@@ 3 Model Specifications
In this section, we provide a number of specific cases in which the proposed vitality-based framework is implemented to arrive at various mortality modeling outcomes. The proposed framework is designed to be modulable, meaning that the four components can be selectively chosen based on different modeling objectives. For instance, by controlling certain components, the framework can replicate traditional mortality laws, offering an alternative approach to basic mortality analysis. Conversely, by incorporating all components, the resulting model is capable of capturing more complex phenomena that influence mortality dynamics over time. Through the specific cases, we aim to demonstrate the capabilities of the vitality-based framework across different modeling needs and highlight its potential to enhance our understanding of mortality patterns.
3.1 Static cases
A vitality mortality model is static if it only governs the mortality dynamics over an individual’s lifetime but does not account for general changes in population mortality over time. We discuss the dynamic cases in the next subsection, where period and cohort effects are captured and reflected by a vitality mortality model.
3.1.1 Traditional mortality laws
Let us consider the following specification of the vitality-based mortality modeling framework:
@@ 3.1.1 Traditional mortality laws
    The initial vitality $V(0)$ follows an exponential distribution with unit scale; that is,
    $V(0)\sim\text{Exp}(1).$
@@ 3.1.2 Plateau death models
    The trend component $Y(t)$ is governed by a mortality law $\mu_{x}(t)$:
    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
@@
An alternative specification for capturing the plateau effect and resulting in equation (3) for the $T$-year survival probability is
@@
    The initial vitality $V(0)$ follows an exponential distribution with unit scale:
    $V(0)\sim\text{Exp}(1).$
@@
    The trend component $Y(t)$ is given by
    where $\mu_{x}(t)$ is specified by a mortality law and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$.
These two specifications have distinct interpretations. The first one describes the initial vitality level of the underlying population to be more dispersed and the reason for the death plateau is due to a larger proportion of the population exhibiting a large initial vitality level. The second one implies that the death plateau is indeed due to the heterogeneity in the depletion rate.
@@ 3.1.3 A jump-diffusion vitality model
To introduce randomness, Shimizu et al., (2021) and Shimizu et al., (2023) have considered applying a Brownian motion and an inverse-Gaussian process, respectively, to govern an individual’s vitality. However, these studies assume a deterministic initial vitality and an arbitrarily chosen trend component, which consequently fails to connect the vitality-based approach with traditional mortality laws. One disadvantage of using inverse Gaussian processes is that their implied infinite number of small jumps may not be suitable for characterizing the decrement of vitality levels due to accidents.
Under the proposed framework, we now consider adding a Brownian motion in the diffusion component to reflect random fluctuations in vitality dynamics, and a compound Poisson process in the jump component to reflect the possibility of accidental events. In particular, the following specification is used:
@@
    The initial vitality $V(0)$ follows a distribution $F_{0}$ with positive support:
    $V(0)\sim F_{0}.$
@@
    The trend component $Y(t)$ is governed by
    $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
    where $\mu_{x}(t)$ is specified by a mortality law (e.g., $\mu_{x}(t)=bc^{x+t}$ for the Gompertz law).
@@
    The diffusion component $W(t)$ is specified as
    $W(t)=\sigma B(t),\quad t\geq 0,$
    where $B(t)$ is a Brownian motion and $\sigma$ is the diffusion parameter.
@@
    The jump component $J(t)$ is specified as
    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables with distribution function $F_{Z}(\cdot)$.
It follows that the dynamics of $V(t)$ is governed by
$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i}.$ (4)
@@
We would like to point out that deriving analytical expressions for the survival probability is generally challenging in this case, unless specific forms of $Y(t)$ are assumed such as linear functions. To elaborate further, the main difficulty lies in deriving the law of the first passage time of $B(t)$ over a moving boundary, which is a non-trivial problem in applied probability (see, e.g., Salminen, (1988)). Furthermore, deriving analytical results about the death time using renewal arguments may also not be applicable if the trend does not preserve the renewal structure of the vitality dynamics, for instance, when $Y(t)$ is modeled under the Gompertz law. In such cases, one may resort to numerical methods to approximate the survival probability of a jump-diffusion vitality model. In this paper, we employed the Monte Carlo method as suggested in Jin and Wang, (2017) to numerically evaluate the survival probability as indicated in Eq. (5), and provided the details in Appendix B.
3.1.4 An alternative specification
In Section 3.1.2, we illustrated two model specifications that equally describe the death plateau. It is not surprising that more than one specification of the vitality-based modeling framework can provide outcomes that align with mortality laws. In this subsection, we formulate an alternative specification that also connects with mortality laws.
The following specification is assumed:
@@ 3.1.4 An alternative specification
    The initial vitality $V(0)$ follows a Gompertz distribution with shape parameter $\eta$ and scale parameter being one. The distribution function of this Gompertz distribution is
    $F(v)=1-e^{-\eta\left(e^{v}-1\right)}.$
@@
    The trend component $Y(t)$ has a linear decay at the rate of $\delta$; that is,
    $\mathop{}\!\mathrm{d}Y(t)=\delta\mathop{}\!\mathrm{d}t,\quad t\geq 0,$
@@
To introduce randomness in the evolution of an individual’s vitality over time, we can additionally assume that
  • C3
    The diffusion component $W(t)$ is specified as
    $W(t)=\sigma B(t),\quad t\geq 0,$
    where $B(t)$ is a Brownian motion and $\sigma>0$ is the diffusion parameter.
  • C4
    The jump component $J(t)$ is specified as
    $J(t)=\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0,$
    where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of i.i.d. positive random variables.
It follows that the dynamics of $V(t)$ is governed by
@@
One special case worth noting is when the jump size $Z_{i}=\infty$ (i.e., each jump represents a fatal event). In this case, we would have
@@
where the last equation follows from the well-known law of the first passage time of a Brownian motion with linear drift (see, e.g., Siegmund, (1986)).
In general, conditioning on $V(0)$, $V(t)$ is a spectrally negative Lévy process (SNLP), and the Laplace transform of $\tau$ is known explicitly. This allows for the use of either analytical or numerical methods of Laplace inversion to obtain the survival probability. To illustrate this, in Appendix A.2, we provide related results for the vitality mortality model described by equation (7) when the jump distribution is a mixture of exponential distributions.
@@
and the individual dies if $\tilde{V}(t)$ cross the threshold $D=1$,
$\tau=\inf\left\{t\geq 0:\tilde{V}(t)\leq 1\right\}.$
@@
3.2 Dynamic cases
A vitality mortality model is called dynamic when it adapts to changes in mortality over time, thereby explaining the period and cohort effects of mortality data. In stochastic mortality modeling, the period effect refers to changes in mortality over time across ages, while the cohort effect refers to changes in mortality across birth years that remain consistent over the cohort’s lifetime. To capture period and cohort effects, time-varying parameters are needed in a vitality mortality model, similar to the structure of the LC model and the CBD model.
Similar to the static cases where we established a linkage between vitality models and mortality laws, in this subsection we aim to build a bridge between the dynamic vitality mortality models and the popular stochastic mortality models. In particular, we begin by examining a dynamic extension of the Gompertz model, which is connected to the CBD model. We then expand our investigation to other mortality models and also consider cohort effects. We demonstrate that the proposed vitality-based framework is capable of providing new insights and enhancements to existing models.
3.2.1 A dynamic vitality model
@@
Since $Y(t)$ is a non-decreasing process, the $T$-year survival probability can be expressed as
$\Pr(\tau>T)$ $=\mathbb{E}\left[\exp\left(-Y(T)\right)\right].$
@@
The $T$-year survival probability of an individual aged $x$ at time 00 under this continuous version of the CBD model is
$\mathbb{E}\left[\exp\left(-\int_{0}^{T}\mu(x+t,t)\mathop{}\!\mathrm{d}t\right)\right],$
where
$\mu(x+t,t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right)\cdot\left[\exp\left(\kappa_{2}(t)\right)\right]^{(x+t)}.$
@@
$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\cdot\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right).$
@@
Our dynamic vitality mortality model can be constructed to have the same survival probability as other stochastic mortality models, whenever the force of mortality is guaranteed to be non-negative. For instance, consider the Lee-Carter model:
@@
If $\alpha(x)$ and $\beta(x)$ are specified as continuous deterministic functions and $\kappa(t)$ is governed by a continuous random walk with drift, then a dynamic vitality mortality model can be established to match the likelihood of the Lee-Carter model. Another example is the Cox-Ingersoll-Ross (CIR) model for modeling the log force of mortality, as discussed in Huang et al., (2022) and Ungolo et al., (2023), where the likelihood aligns with our vitality model if we maintain the same specification for the depletion rate $\mu_{x}(t)$.
@@
3.2.2 A dynamic vitality model with cohort effect
In a factor-based mortality model, the cohort effect is captured by an additive term, often denoted by $\gamma$, such as in the Renshaw-Haberman model (M2), the Age-Period-Cohort model (M3), the CBD model with cohort effect (M6), the CBD model with quadratic and cohort effects (M7), and the CBD model with age-dependent cohort effect (M8), as outlined in Cairns et al., (2009). In a vitality-based mortality model, the cohort effect can be naturally embedded in the initial vitality to reflect changes in the average healthiness of each generation.
We illustrate the idea of a dynamic vitality model with cohort effect based on the dynamic Gompertz law with cohort effect, which corresponds to the CBD model with cohort effect (M6). Denote $V_{y}(t)$ as the time-$t$ vitality level of an individual born in year $y$ with an age of $x$ at time 00. The following specification is assumed:
@@ 3.2.2 A dynamic vitality model with cohort effect
    The trend component $Y(t)$ is specified as in the dynamic Gompertz model:
    $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t,\quad t\geq 0.$
    The log transforms of $c(t)$ and $b(t)$ are governed by
@@
where $\tau_{y}$ is the first passage time for $V_{y}(t)$ crossing zero.
The cohort effect is reflected by the rate parameter $\Gamma(y)$, which can be intuitively interpreted as the average initial vitality level at the time of modeling for individuals born in birth year $y$. The initial condition $\Gamma(0)=1$ ensures identifiability between $\Gamma(y)$ and $b(t)$. Note that the dynamic of $\Gamma(y)$ can be chosen to satisfy certain properties, for example, setting $\mu_{\gamma}=0$ ensures that the future cohorts are expected to be the same as the current one, or modeling $\Gamma(y)$ as a CIR process to incorporate a mean-reverting cohort effect.
We now establish a link between this dynamic vitality model and the continuous version of the M6 model. We fix age $x$ and construct the M6 model with time $t$ and cohort $y$ as
@@
$b(t)=\exp\left(\kappa_{1}(t))-\kappa_{2}(t)\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right),\quad\Gamma(y)=\exp(\gamma(y)).$
@@
3.2.3 A dynamic vitality model with age-cohort effect
The dynamic vitality models presented in Sections 3.2.1 and 3.2.2 begin the modeling of vitality with a fixed initial age. It would be interesting to formulate a vitality model that allows the vitality to start from an arbitrary age. We now consider a dynamic vitality model with an age-adjusted cohort effect, extended again from the Gompertz Law.
The following specification is applied:

### v3

@@ 3 Model Specifications
In this section, we present several vitality models to illustrate how the proposed framework can be tailored to achieve different mortality modeling outcomes and also to enhance existing mortality models. Figure 1 outlines how the remainder of this section is structured and also highlights the connection between our vitality models and some existing models.
Vitality Mortality ModelStatic ParametersDynamic Parameters Plateau Models Mortality Laws Jump-Diffusion Models M5 Model M6 Model M8 Model Section 3.1.1Section 3.1.2Section 3.1.3Section 3.2.1Section 3.2.2Section 3.2.3
Figure 1: Connections between vitality mortality models and existing mortality models.
3.1 Models with static parameters
We begin from vitality models that are built by static parameters. The first goal is to establish a connection between traditional mortality laws and vitality models. We then extend those vitality models to include diffusion and jump components and derive analytical solutions for the resulting survival probabilities.
3.1.1 Mortality laws
Let us consider the following specification of the proposed framework:
@@ 3.1.2 Plateau death models
    The trend component $Y(t)$ is governed by a mortality law $\mu_{x}(t)$: $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t$.
@@
An alternative specification for capturing the plateau effect is
@@
    The initial vitality follows an exponential distribution with unit scale: $V(0)\sim\text{Exp}(1)$.
@@
    The trend component is given by $\mathop{}\!\mathrm{d}Y(t)=Z\mu_{x}(t)\mathop{}\!\mathrm{d}t$, where $\mu_{x}(t)$ is specified by a mortality law and $Z$ follows a Gamma distribution with shape parameter $\alpha$ and rate parameter $\phi$.
The resulting $T$-year survival probability of the above specification is the same as equation (3) (see Appendix A.1 for more details). These two specifications have distinct interpretations. The first one describes the initial vitality level of the underlying population as being more dispersed and the reason for the death plateau is due to a larger proportion of the population exhibiting a large initial vitality level. The second one implies that the death plateau is indeed due to the heterogeneity in the depletion rate. While it is possible to find other formulations that yield the same survival probabilities, our study does not aim to be exhaustive in covering all possible formulations. We provide an additional formulation at the end of Appendix A.1.
@@ 3.1.3 A jump-diffusion vitality model
To introduce randomness, Shimizu et al., (2021) and Shimizu et al., (2023) have considered applying a diffusion process and an inverse-Gaussian process, respectively, to govern an individual’s vitality. However, these studies assume a deterministic initial vitality and an arbitrarily chosen trend component, which consequently fails to connect the vitality-based approach with traditional mortality laws. Moreover, one disadvantage of using inverse Gaussian processes is that their implied infinite number of small jumps may not be suitable for characterizing the decrement of vitality levels due to accidents.
Under the proposed framework, we now consider adding a Brownian motion in the diffusion component to reflect random fluctuations in vitality dynamics, and a compound Poisson process in the jump component to reflect the possibility of accidental events. In particular, a jump-diffusion vitality mortality model is specified as follows:
@@
    The initial vitality $V(0)$ follows a distribution $F_{0}$ with positive support: $V(0)\sim F_{0}$.
@@
    The trend component $Y(t)$ is governed by $\mathop{}\!\mathrm{d}Y(t)=\mu_{x}(t)\mathop{}\!\mathrm{d}t$, where $\mu_{x}(t)$ is specified by a mortality law.
@@
    The diffusion component $W(t)$ is specified as $W(t)=\sigma B(t)$, where $B(t)$ is a Brownian motion and $\sigma$ is the diffusion parameter.
@@
    The jump component $J(t)$ is specified as $J(t)=\sum_{i=1}^{N(t)}Z_{i}$, where $N(t)$ is a Poisson process with intensity rate $\lambda(t)$, and $\{Z_{i}\}_{i\geq 1}$ is a sequence of independent and identically distributed (i.i.d.) positive random variables with distribution function $F_{Z}(\cdot)$.
The dynamics of $V(t)$ is consequently governed by
$V(t)=V(0)-\int_{0}^{t}\mu_{x}(s)\mathop{}\!\mathrm{d}s-\sigma B(t)-\sum_{i=1}^{N(t)}Z_{i},\quad t\geq 0.$ (4)
@@
We remark that deriving analytical expressions for the survival probability is generally challenging in this case, unless specific forms of $Y(t)$ are assumed such as linear functions. To elaborate further, the main difficulty lies in deriving the law of the first passage time of $B(t)$ over a moving boundary, which is a non-trivial problem in applied probability (see, e.g., Salminen, (1988)). Furthermore, deriving analytical results about the death time using renewal arguments may also not be applicable if the trend does not preserve the renewal structure of the vitality dynamics, for instance, when $Y(t)$ is modeled under the Gompertz law. In such cases, one may resort to numerical methods to approximate the survival probability of a jump-diffusion vitality model. In this paper, we employed the Monte Carlo method as suggested in Jin and Wang, (2017) to numerically evaluate the survival probability as indicated in equation (5), and provided the details in Appendix B.
In Section 3.1.2, we illustrated two model specifications that equally describe the death plateau. It is not surprising that more than one specification of the vitality-based modeling framework can provide outcomes that align with mortality laws. We now formulate an alternative specification that also connects with mortality laws. The following specification is assumed:
@@
    The initial vitality $V(0)$ follows a Gompertz distribution with shape parameter $\eta$ and scale parameter being one. The distribution function of this Gompertz distribution is $F(v)=1-e^{-\eta\left(e^{v}-1\right)}.$
@@
    The trend component $Y(t)$ has a linear decay at the rate of $\delta$; that is, $\mathop{}\!\mathrm{d}Y(t)=\delta\mathop{}\!\mathrm{d}t$.
@@
To introduce randomness in the evolution of an individual’s vitality over time, we can add the diffusion component and the jump component that were previously assumed in this subsection to formulate the jump-diffusion vitality mortality model again. The dynamics of $V(t)$ will be governed by
@@
One special case worth noting is when the jump represents a fatal event and the survival probability can be analytically expressed in the following proposition.
Proposition 1.
Given the vitality model (7) with $Z_{i}=\infty$, the survival probability can be analytically solved as
@@
where $\Phi(\cdot)$ is the distribution function of a standard Gaussian distribution.
Proposition 1 directly follows from the law of the first passage time of a Brownian motion with linear drift (see, e.g., Siegmund, (1986)), and it is thus omitted here. In general, conditioning on $V(0)$, $V(t)$ is a spectrally negative Lévy process (SNLP), and the Laplace transform of $\tau$ is known explicitly. This allows for the use of either analytical or numerical methods of Laplace inversion to obtain the survival probability. To illustrate this, in Appendix A.2, we provide related results for the vitality mortality model described by equation (7) when the jump distribution is a mixture of exponential distributions.
@@
3.2 Models with dynamic parameters
In this subsection, we build vitality models that use on dynamic parameters and consequently have a connection with stochastic mortality models. We begin by examining a dynamic extension of the Gompertz law, and then expand our investigation to models with age and cohort effects.
3.2.1 A vitality model with period effect
@@
    The initial vitality $V(0)$ follows an exponential distribution with unit scale: $V(0)\sim\text{Exp}(1)$.
@@
Since $Y(t)$ is a non-decreasing process, the $T$-year survival probability can be expressed as $\Pr(\tau>T)=\mathbb{E}\left[\exp\left(-Y(T)\right)\right]$.
@@
The $T$-year survival probability of an individual aged $x$ at time 00 under this continuous version of the CBD model is $\mathbb{E}\left[\exp\left(-\int_{0}^{T}\mu(x+t,t)\mathop{}\!\mathrm{d}t\right)\right]$, where
$\mu(x+t,t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\cdot\bar{x}\right)\cdot\left[\exp\left(\kappa_{2}(t)\right)\right]^{x+t}.$
@@
$b(t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\cdot\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right).$
@@
Our vitality mortality model can be constructed to have the same survival probability as other stochastic mortality models, whenever the force of mortality is guaranteed to be non-negative. For instance, consider the Lee-Carter model:
@@
If $\alpha(x)$ and $\beta(x)$ are specified as continuous deterministic functions and $\kappa(t)$ is governed by a continuous random walk with drift, then a vitality mortality model can be established to match the likelihood of the Lee-Carter model. Another example is the Cox-Ingersoll-Ross (CIR) model for modeling the log force of mortality, as discussed in Huang et al., (2022) and Ungolo et al., (2024), where the likelihood aligns with our vitality model if we maintain the same specification for the depletion rate $\mu_{x}(t)$.
@@
3.2.2 A vitality model with period-cohort effect
In a factor-based mortality model, the cohort effect is captured by an additive term, often denoted by $\gamma$, such as in the Renshaw-Haberman model (M2), the Age-Period-Cohort model (M3), the CBD model with cohort effect (M6), the CBD model with quadratic and cohort effects (M7), and the CBD model with age-dependent cohort effect (M8), as outlined in Cairns et al., (2009). In a vitality mortality model, the cohort effect can be naturally embedded in the initial vitality to reflect changes in the average healthiness of each generation.
We illustrate the idea of a vitality model with cohort effect based on the dynamic Gompertz law with cohort effect, which corresponds to the CBD model with cohort effect (M6). Denote $V_{y}(t)$ as the time-$t$ vitality level of an individual born in year $y$ with an age of $x$ at time 00. The following specification is assumed:
@@ 3.2.2 A vitality model with period-cohort effect
    The trend component $Y(t)$ is specified as in the dynamic Gompertz model: $\mathop{}\!\mathrm{d}Y(t)=b(t)c(t)^{x+t}\mathop{}\!\mathrm{d}t$, $t\geq 0$. The log transforms of $c(t)$ and $b(t)$ are governed by
@@
where $\tau_{y}$ is the first passage time for $V_{y}(t)$ crossing zero, and one can refer to Appendix A.3 for more details.
The cohort effect is reflected by the rate parameter $\Gamma(y)$, which can be intuitively interpreted as the average initial vitality level at the time of modeling for individuals born in birth year $y$. The initial condition $\Gamma(0)=1$ ensures identifiability between $\Gamma(y)$ and $b(t)$. Note that the dynamics of $\Gamma(y)$ can be chosen to satisfy certain properties, for example, setting $\mu_{\gamma}=0$ ensures that the future cohorts are expected to be the same as the current one, or modeling $\Gamma(y)$ as a CIR process to incorporate a mean-reverting cohort effect.
We now establish a link between this vitality mortality model and the continuous version of the M6 model. We fix age $x$ and construct the M6 model with time $t$ and cohort $y$ as
@@
$b(t)=\exp\left(\kappa_{1}(t)-\kappa_{2}(t)\bar{x}\right),\quad c(t)=\exp\left(\kappa_{2}(t)\right),\quad\Gamma(y)=\exp(\gamma(y)).$
@@
3.2.3 A vitality model with age-adjusted period-cohort effect
The vitality models presented in Sections 3.2.1 and 3.2.2 begin the modeling of vitality with a fixed initial age $x$. It would be interesting to formulate a vitality model that allows the vitality to start from an arbitrary age and the cohort effect to vary as the individual ages. To achieve this, we consider a vitality model with an age-adjusted cohort effect, extended again from the Gompertz Law. The following specification is applied: